# Shaving Weights with Occam’s Razor: Bayesian Sparsification for Neural Networks using the Marginal Likelihood

### v1

@@ B.8 Network Compression
In Figures 18 and 19, we demonstrate the efficiency gains achieved by our SpaM-OPD approach. For the fully connected network on the Cancer dataset, it achieves a remarkable reduction of over 20 times in disk size and 24 times in FLOPs while simultaneously maintaining baseline test accuracy. Additionally, it boasts a Brier score of 0.15 and a negative log marginal likelihood (Neg Log MargLik) lower than the original model. These results highlight the effectiveness of SpaM-OPD in achieving significant model compression without compromising performance on key metrics.
@@
Figure 18: Structured and unstructured pruning of LeNet on FashionMNIST with SpaM-OPD. We see that through structured sparsification, we are able to obtain models that are still performant at a significantly reduced computational and memory cost, while unstructured pruning does not directly translate into computational benefits.
@@
Figure 19: Structured and unstructured pruning of FC on Cancer with SpaM-OPD. We see that through structured sparsification, we are able to obtain models that are still performant at a significantly reduced computational and memory cost. At the same time, unstructured pruning does not directly translate into computational benefits.
@@ D.1 Datasets
CANCER (UCI): This dataset, derived from digitized images of fine needle aspirates of breast masses, includes features describing characteristics of cell nuclei in the images. It was pivotal for breast cancer research (Dua and Graff, 2019).
MNIST: A foundational benchmark dataset in machine learning, MNIST consists of handwritten digits (0 to 9) in 28x28 pixel grayscale images (LeCun et al., 1998).
FashionMNIST: A more challenging variant of MNIST, FashionMNIST encompasses 60,000 training and 10,000 test images in grayscale (28x28 pixels), each representing one of ten clothing categories (Xiao et al., 2017).
CIFAR-10: 60,000 colored images (32x32 pixels) divided equally among ten distinct classes (Krizhevsky, 2009). We augment the dataset using random flipping and cropping for ResNet in the cases where we explicitly mention it.
D.2 Neural Network Models
@@
LeNet: As a foundational Convolutional Neural Network (CNN), LeNet has shown exceptional performance in digit and image recognition tasks. We have applied LeNet to the MNIST, Fashion MNIST, and CIFAR-10 datasets, leveraging its capability to handle varying complexities of image data (LeCun et al., 1998). LeNet on CIFAR-10 is not a very common benchmark for pruning; here, it is used to demonstrate how SpaM, and specifically SpaM-OPD, is able to prune at high percentages without a performance loss up to 80% in a model that struggles with representing the data’s complexity, showing that our work extends beyond over-parametrized networks for the task at hand.
@@
MLPMixer: The MLPMixer serves as a streamlined alternative to more complex models like CNNs and transformers. It relies solely on Multi-Layer Perceptrons (MLPs) for integrating inputs across spatial and channel dimensions (Tolstikhin et al., 2021). We implement an MLPMixer with 2 blocks designed for MNIST.
@@
ResNet with inplanes 64 and depth 18 for CIFAR-10: We modify the implementation of ResNet and incorporate fixup initialization and custom bias and scale parameters to align with the constraints of the ASDL backend (Osawa et al., 2023) used for the Laplace computations in this work, which does not support batch normalization.
@@
Wide ResNet: decreases depth compared to ResNet and increases the width of residual networks (Zagoruyko and Komodakis, 2017) with a depth of 16 and a widening factor of 4 (WRN16-4). We use fixup blocks to be able to utilize ASDL backend (Osawa et al., 2023).
@@
Vision Transformer (ViT) (Dosovitskiy et al., 2021): unlike CNNs, which extract local features through filters and pooling layers, ViT breaks down images into fixed-size patches, treating each as a "token" in a sequence (Dosovitskiy et al., 2021). This allows it to leverage the Transformer architecture, initially designed for language processing, to analyze relationships between patches through self-attention mechanisms (Vaswani et al., 2017).
D.3 Pruning Criteria
@@
  • SNIP: Uses connection sensitivity, how much a specific weight contributes to the output loss, for effective pruning (Lee et al., 2018).
@@
    $GF=gL(\Theta)^{T}gL(\Theta)=||gL(\Theta)||^{2}_{2},$ (9)
    emphasizing the impact of pruning on the training dynamics (Wang et al., 2020). We replicate the GraSP implementation of Rachwan et al. (2022), where we consider the absolute value of the importance score initially proposed by Lubana and Dick (2021) given by:
    $I(\Theta_{t})=|\Theta_{t}^{T}H_{L}(\Theta_{t})g_{L}(\Theta_{t})|$ (10)
@@
  • Structured-SynFLow: We challenge the capabilities of SynFlow (Tanaka et al., 2020), a data-agnostic pruning approach that prevents layer collapse that happens at high sparsities where layers are no longer able to perform at the model’s predictive power. This typically occurs when the pruning algorithm, intentionally or inadvertently, removes a significant portion of weights or filters from a specific layer, effectively collapsing its functionality (Tanaka et al., 2020). We push SynFlow to its limits through advanced structured pruning strategies, where we prune layers aggressively at the same target sparsity, which facilitates the compression process and resizing. By applying rigorous layer-specific filtering and neuron pruning, we aim to test the robustness and effectiveness of SynFlow in extreme sparsity structured scenarios. This approach not only benchmarks SynFlow’s performance under stringent conditions but also explores its potential to maintain network functionality and accuracy in highly sparse neural network architectures.
  • Magnitude Pruning: Relies on the magnitude of weights for pruning, aiming to maintain model performance while reducing complexity (Han et al., 2015). After the success shown by Han et al. (2015), many methods adapted magnitude as a pruning criterion coupled with different scheduling (Bellec et al., 2018; Mostafa and Wang, 2019; Zhou et al., 2020).
@@
D.4 Structured Sparsification Process

### v2

@@ B.10 Network Compression
In Figures B11 and B12, we demonstrate the efficiency gains achieved by our SpaM-OPD approach. For the fully connected network on the Cancer dataset, it achieves a remarkable reduction of over 20 times in disk size and 24 times in FLOPs while simultaneously maintaining baseline test accuracy. Additionally, it boasts a Brier score of 0.15 and a negative log marginal likelihood (Neg Log MargLik) lower than the original model. These results highlight the effectiveness of SpaM-OPD in achieving significant model compression without compromising performance on key metrics.
@@
Figure B11: Structured and unstructured pruning of LeNet on FashionMNIST with SpaM-OPD. We see that through structured sparsification, we are able to obtain models that are still performant at a significantly reduced computational and memory cost, while unstructured pruning does not directly translate into computational benefits.
@@
Figure B12: Structured and unstructured pruning of FC on Cancer with SpaM-OPD. We see that through structured sparsification, we are able to obtain models that are still performant at a significantly reduced computational and memory cost. At the same time, unstructured pruning does not directly translate into computational benefits.
@@ D.1 Datasets
  • Breast Cancer Wisconsin (Diagnostic) (UCI): This dataset, derived from digitized images of fine needle aspirates of breast masses, includes features describing characteristics of cell nuclei in the images. It is a classic binary classification dataset used extensively in breast cancer research [57].
  • MNIST: A foundational benchmark dataset in machine learning, MNIST consists of 60,000 training and 10,000 test images of handwritten digits (0 to 9) in 28x28 pixel grayscale format [58].
  • FashionMNIST: A drop-in replacement for MNIST, Fashion-MNIST offers a greater challenge with its 60,000 training and 10,000 test images in grayscale (28x28 pixels). Each image represents one of ten clothing categories [59].
  • CIFAR-10: This dataset contains 60,000 color images (32x32 pixels) divided equally among 10 classes (e.g., airplane, bird, cat) [60]. For our ResNet experiments, we augment CIFAR-10 with random flipping and cropping.
  • CIFAR-100: A more fine-grained version of CIFAR-10, this dataset includes 60,000 color images (32x32 pixels) across 100 classes, with 600 images per class [60]. We apply random flipping and cropping for augmentation.
  • IMDB Movie Review: This dataset is a collection of 50,000 movie reviews, balanced between positive and negative sentiments. It is commonly used for binary sentiment classification tasks [61].
D.2 Models
@@
  • LeNet: As a foundational Convolutional Neural Network (CNN), LeNet has shown exceptional performance in digit and image recognition tasks. We have applied LeNet to the MNIST, Fashion MNIST, and CIFAR-10 datasets, leveraging its capability to handle varying complexities of image data [58]. LeNet on CIFAR-10 is not a very common benchmark for pruning; here, it is used to demonstrate how SpaM, and specifically SpaM-OPD, is able to prune at high percentages without a performance loss up to 80% in a model that struggles with representing the data’s complexity, showing that our work extends beyond over-parametrized networks for the task at hand.
@@
  • MLPMixer: The MLPMixer serves as a streamlined alternative to more complex models like CNNs and transformers. It relies solely on Multi-Layer Perceptrons (MLPs) for integrating inputs across spatial and channel dimensions [62]. We implement an MLPMixer with 2 blocks designed for MNIST.
@@
  • ResNet with inplanes 64 and depth 18 for CIFAR-10: We modify the implementation of ResNet and incorporate fixup initialization and custom bias and scale parameters to align with the constraints of the ASDL backend [63] used for the Laplace computations in this work, which does not support batch normalization.
@@
  • Wide ResNet: decreases depth compared to ResNet and increases the width of residual networks [64] with a depth of 16 and a widening factor of 4 (WRN16-4). We use fixup blocks to be able to utilize ASDL backend [63].
@@
  • Vision Transformer (ViT) [65]: unlike CNNs, which extract local features through filters and pooling layers, ViT breaks down images into fixed-size patches, treating each as a "token" in a sequence [65]. This allows it to leverage the Transformer architecture, initially designed for language processing, to analyze relationships between patches through self-attention mechanisms [66].
  • DistilBERT: DistilBERT [67] is a smaller, faster, and cheaper version of BERT, achieved by leveraging knowledge distillation during the pre-training phase. This model retains 97% of BERT’s language understanding capabilities while being 60% faster and 40% smaller. We use the pre-trained DistilBERT hosted in Hugging Face under (distilbert-base-uncased) [67] and tune it for sentiment analysis to classify reviews in the IMDB dataset [61] , which involves predicting the sentiment (positive or negative) of user reviews based on their textual content.
  • GPT-2: a large-scale transformer-based language model developed by OpenAI, with impressive text generation capabilities. Trained on a vast corpus of internet text [68]. In our study, we leverage the 124M parameter version of GPT-2, fine-tuning it on the IMDB dataset for sentiment analysis to assess its performance under different pruning conditions.
D.3 Dependencies
For the computation of second-order information (e.g., Hessian, Fisher information) needed for the Laplace approximation, we utilize the ASDL Library [63]. We use the library in its version under https://github.com/kazukiosawa/asdl/tree/011a942b2698b9ec33b0c8c47c96bd49335e5d80. The ASDL Library is distributed under the MIT License, which allows for reuse with a few restrictions that we respect in our work.
D.4 Hyperparameters
Marginal Likelihood
  • Hessian Approximation: The choice between GGN and EF. GGN was initially employed for fully connected networks, LeNets, and ResNets. However, for complex architectures (WRNs, ViTs, DistilBERT), GGN’s computational cost became prohibitive, exceeding MAP runtime by up to 20x and even more for casual modeling tasks. Switching to EF maintained pruning performance while closely matching MAP runtime, which is particularly beneficial as GGN scales linearly with the number of classes. We discuss further the cost in Section D.7.
  • n_epochs_burnin Dictates the number of epochs after which marginal likelihood optimization starts. If set superior to the number of training epochs, marginal likelihood is skipped, and the training is equivalent to MAP.
  • marglik_frequency Controls the frequency of marginal likelihood estimation. The default value of 1 signifies re-estimation after each epoch, while a value of 5 indicates approximation for every fifth epoch.
We use these parameters to manage the computational cost of our experiments, where for small models like LeNets, FC Networks, the n_epochs_burnin is set to zero and marglik_frequency to one reflecting estimating each epoch since the start of the training. In contrast, for complex networks like MLPMixer, ResNets, WideResNet, and ViT that we train from scratch, we start after 20 epochs and at an interval frequency of 5 epochs.
Hyperparameters Table D4 presents the specific hyperparameters employed for each dataset-architecture combination. We use $\dagger$ to denote the use of data augmentation in the training process. The symbols $\star$ and $\diamond$ represent the use of the Generalized Gauss-Newton (GGN) and Empirical Fisher (EF) approximations for the Hessian, respectively. We use cosine decay scheduler towards a fixed minimum learning rate of 1e-6 across all experiments. The symbols $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, etc., represent the following datasets:
  • $\mathcal{D}_{1}$: Breast Cancer Wisconsin (Diagnostic)
  • $\mathcal{D}_{2}$: MNIST
  • $\mathcal{D}_{3}$: FashionMNIST
  • $\mathcal{D}_{4}$: CIFAR-10
  • $\mathcal{D}_{5}$: CIFAR-100
  • $\mathcal{D}_{6}$: IMDB Movie Review
[…]
Table D4: Hyperparameters used in the experiments.
Dataset (Arch.) Marglik Freq. Batch Size Learning Rate Optimizer Temp. Burn-in / Epochs
$\mathcal{D}_{1}^{\blacktriangle}$ (FCN) 1 64 0.001 Adam 1.0 0 / 50
$\mathcal{D}_{2}^{\blacktriangle}$ (FCN) 1 64 0.001 Adam 1.0 0 / 100
$\mathcal{D}_{2}^{\blacktriangle}$ (LeNet) 1 128 0.001 SGD 1.0 0 / 100
$\mathcal{D}_{3}^{\blacktriangle}$ (LeNet) 1 128 0.001 SGD 1.0 0 / 100
$\mathcal{D}_{3}^{\blacktriangle}$ (MLPMixer) 1 128 0.001 Adam 1.0 0 / 100
$\mathcal{D}_{4}^{\dagger}$ (ResNet) 5 128 0.1 SGD 5 20 / 100
$\mathcal{D}_{5}^{\dagger}$ (WRN) 5 128 0.1 SGD 5 20 / 200
$\mathcal{D}_{2}^{\blacktriangle}$ (ViT) 5 128 0.001 Adam 1.0 20 / 100
$\mathcal{D}_{6}^{\blacktriangledown}$ (DistilBERT) 5 32 2e-5 AdamW 1.0 5 / 20
$\mathcal{D}_{6}^{\blacktriangledown}$ (GPT-2) 5 8 2e-5 Adam 1.0 5 / 10
Computational resources Our experiments are run on GPUs. We run our experiments in a single GPU configuration on available variation between 1080 Tis, V100s, and A100s, with the majority being run on A100s with 40GB memory as we run the experiments intensively one after the other for different architecture on the same allocated GPU and in order to provide enough GPU memory. For models such as FCs, LeNets, ResNets, and MLP-Mixer, a GPU with 12GB of memory ( 1080 Ti) proved sufficient to run our method for our recommended laplace and prior, which is diagonal with parameter-wise priors and reproduce the results. For the sentiment analysis task using GPT-2, we recommend using a 32 GB GPU for tuning to be able to utilize a high batch size and to use diagonal approximation to fit laplace on the data without running into memory shortage.
Runtime
Table D5 presents the training and pruning runtimes on A100 for each dataset-architecture combination. Training times are given for both SpaM diagonal with parameter-wise prior and MAP, while pruning time is identical to both. Pruning runtimes refer to the time taken for OPD to compute and prune a model at 10 target sparsities. OPD and magnitude are very close in terms of runtime and the most efficient compared to SNIP, which is slightly slower due to it requiring an additional forward pass, and GraSP, which is significantly slower as it accumulates the gradient as shown in Figure D13.
Table D5: Runtimes for the experiments. Train: Training time (h:m), Prune: Pruning time (m:s).
Dataset (Arch.) Train Prune
SpaM MAP OPD
$\mathcal{D}_{1}^{\blacktriangle}$ (FCN) 0:01 0:01 0:04
$\mathcal{D}_{2}^{\blacktriangle}$ (FCN) 0:15 0:5 0:23
$\mathcal{D}_{2}^{\blacktriangle}$ (LeNet) 0:16 0:10 0:15
$\mathcal{D}_{3}^{\blacktriangle}$ (LeNet) 0:16 0:10 0:21
$\mathcal{D}_{3}^{\blacktriangle}$ (MLPMixer) 0:05 0:07 0:10
$\mathcal{D}_{4}^{\dagger}$ (ResNet) 1:24 0:25 0:55
$\mathcal{D}_{5}^{\dagger}$ (WRN) 1:17 1:12 0:51
$\mathcal{D}_{2}^{\blacktriangle}$ (ViT) 0:26 0:15 0:24
$\mathcal{D}_{6}^{\blacktriangledown}$ (DistilBERT) 5:20 2:15 1:05
$\mathcal{D}_{6}^{\blacktriangledown}$ (GPT-2) 17:34 6:24 17:41
Figure D13: Mean relative pruning time with standard error bars on WRN with CIFAR-100. OPD and Magnitude are the most efficient as they use pre-computed parameters, with SNIP being slower due to requiring an additional forward pass, while GraSP is significantly slower as it needs to accumulate the gradient.
D.5 Pruning Criteria
@@
  • SNIP: Uses connection sensitivity, how much a specific weight contributes to the output loss, for effective pruning [36].
@@
    $GF=gL(\Theta)^{T}gL(\Theta)=||gL(\Theta)||^{2}_{2},$ (D1)
    emphasizing the impact of pruning on the training dynamics [37]. We replicate the GraSP implementation of Rachwan et al. [39], where we consider the absolute value of the importance score initially proposed by Lubana and Dick [38] given by:
    $I(\Theta_{t})=|\Theta_{t}^{T}H_{L}(\Theta_{t})g_{L}(\Theta_{t})|$ (D2)
@@
  • Structured-SynFLow: We challenge the capabilities of SynFlow [50], a data-agnostic pruning approach that prevents layer collapse that happens at high sparsities where layers are no longer able to perform at the model’s predictive power. This typically occurs when the pruning algorithm, intentionally or inadvertently, removes a significant portion of weights or filters from a specific layer, effectively collapsing its functionality [50]. We push SynFlow to its limits through advanced structured pruning strategies, where we prune layers aggressively at the same target sparsity, which facilitates the compression process and resizing. By applying rigorous layer-specific filtering and neuron pruning, we aim to test the robustness and effectiveness of SynFlow in extreme sparsity structured scenarios. This approach not only benchmarks SynFlow’s performance under stringent conditions but also explores its potential to maintain network functionality and accuracy in highly sparse neural network architectures.
  • Magnitude Pruning: Relies on the magnitude of weights for pruning, aiming to maintain model performance while reducing complexity [69]. After the success shown by Han et al. [69], many methods adapted magnitude as a pruning criterion coupled with different scheduling [49, 70, 48].
@@
D.6 Structured Sparsification Process
@@
D.7 Computational Cost
Instead of using the Generalized Gauss-Newton (GGN) approximation, which scales linearly with the number of classes, we can also use the Empirical Fisher (EF). For most architectures, using EF instead of GGN for SpaM does not add a very large computational overhead to MAP, as EF costs roughly as much as gradient computation. This is particularly beneficial as GGN scales linearly with the number of classes. The pruning results are not significantly affected by the choice of GGN or EF.
The runtime of MAP and SpaM was close (roughly 1h and 20 minutes on A100s) for WRN-16 on CIFAR100 using SpaM (EF) with diagonal LA and parameter-wise priors (our recommended settings for pruning). For language transformers, specifically DistilBERT and GPT-2, SpaM with EF does result in a longer training time compared to MAP. However, this increase is considerably less than when using GGN, where a single epoch can take longer than the entire SpaM training with EF.
In prior works [71], it was found that GGN gives a better posterior predictive approximation, but we do not use it in this work. We find that EF works similarly well for pruning at a much lower cost.